Mr. Thomas is putting a fence around his cornfield. The field is 360 yards long and 225 yards wide. He wants to put a fence post every 3yards. How many fence posts Will mr.Thomas need?
Found 2 solutions by harpazo, Alan3354:
Answer by harpazo(655)

Let's see.
360/3 = 120
225/3 = 75
P = 2 times width plus 2 times length.
P = 2(75) + 2(120)
P = 150 + 240
P = 390
He will need 390 fences.

Answer by Alan3354(69443)

Mr. Thomas is putting a fence around his cornfield. The field is 360 yards long and 225 yards wide. He wants to put a fence post every 3yards. How many fence posts Will mr.Thomas need?
-------------
On the long sides, he'll need 360/3 + 1 = 121 posts
On the short sides, 225/3 + 1 = 76 would be needed, but the corners already have posts --> 74 each
-----
121*2 + 74*2 = 390 posts